what is the percentage composition of Na2CO3


1
Expert's answer
2021-10-10T10:20:13-0400

Given Compound = Sodium Carbonate (Na2CO3)

Molecular Mass of Na = 23

Molecular Mass of C = 12

Molecular Mass of O = 16

So, molecular mass of Na2CO3 = 2(23) + 12 + 3(16) = 106

Therefore, percentage composition of each element in Sodium Carbonate will be calculated by formula : (Amount of element/Total Molecular Mass)*100%


Na = ((2*23)/106)*100% = 43.39%

C = (12/106)*100% = 11.32%

O = ((3*16)/106)*100% = 45.29%